OpenStudy (anonymous):

Please help will give medal. Solve the following equation for t. v2=v1 + at t= ??

OpenStudy (yolomcswagginsggg):

@barricade1001 look dat up

OpenStudy (anonymous):

v2-v1=at \[\frac{ v2-v1 }{ a }=t\]

OpenStudy (anonymous):

That is the same answer i got but it says it is incorrect.

OpenStudy (anonymous):

see if they have given something else otherwise it is correct.